(2 pts.) For the following electrochemical cell at 25°C: Zn(s) | Zn2+ (aq) || H+ (aq)...

  1. (2 pts.) For the following electrochemical cell at 25°C:

Zn(s) | Zn2+ (aq) || H+ (aq) | Pt,H2 (g) (P=1.00 atm)

  1. calculate DG° in kJ
  2. calculate Ecell for [H+] = 0.8 M and [Zn2+] = 0.5 M
  3. calculate K at 25°C
Answer #1

Zn(s) | Zn​​​​​ 2+ (aq) || H + (aq) | pt , H 2 (g) ( P= 1.00 atm )

For Zn | Zn 2+ standard electrode potential is represent as

E0​​​​​ Zn | Zn​​​​​2+ = 0.76 V and for hydrogen , E​​​​​​0 H2 | 2H​​​​​+ = 0.00 V

L. H .E. Zn (s) Zn​​​​​​​​2+ (aq)  + 2e  

R.H.E. 2H + (aq ) +2e   H​​​​​​2 (g)

Hence the cell reaction is  

Zn(s) + 2H+ (aq) Zn​​​​​2+ (aq) + H​​​​​​2 (g) ............ (1)

Hence the two electron take part in this reaction . Hence n= 2 ,

Hence the total E 0 cell is = E0​​​​​ Zn | Zn​​​​​2+ + E​​​​​​0 2H+ | H2

=( 0.76 + 0.0 ) V

= 0.76 V

a) We know that ∆G0 = - nFE​​​​0

We know that F = 96500C

    ∆G0 = -( 2 mol × 96500 C mol​​​​-1 × 0.76 V )

∆G0 = - 146680 CV = -146680 J

∆G0 = -146.68 KJ

b) From nernst equation we get (using equation 1)

For solid [Zn] = 1

  

Here, For solid [Zn] = 1 [Zn​​​​​2+] = 0.5 M

[H+] = 0.8 M  

E = 0.756 V  

Hence the E cell = 0.756 V

C) At equilibrium E​​​​​​cell = 0

E​​​​​​0​​​Cell = (0.059/n) logK

0.76 = (0.059/2) logK

logK = 25.7627

K = 5.79 × 1025

Hence K = 5.79 × 1025

Here T=250 C = (25+273) K = 298 K

R = 8.314 J mol​​​​-1 K​​​​​​-1